OpenStudy (anonymous):

ANSWER WITH ALL THE STEPS OF WORK GETS A MEDAL! Write the equation of the polynomial function f(x) that has roots x=2, x=1 and x=3 and has a y-intercept of (0,24)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

start with \[f(x)=(x-2)(x-1)(x-3)\] you need to find \(a\) olve for \(a\) by putting in a \(0\) for \(x\) , setting the result equal to \(24\) and solving for \(a\)

OpenStudy (anonymous):

f(x)=(0−2)(0−1)(0−3) f(24)=(0−2)(0−1)(0−3) and what do i do from here @misty1212 ???